OpenStudy (theqube1):

Geometry help Connections Academy Area Unit Test Main Lesson Content. Geometry B Lesson 9 Unit 4

OpenStudy (theqube1):

Help Me Please @BugzytheGreat

OpenStudy (theqube1):

If you have the full answers that works since the test is only 16 questions

OpenStudy (anonymous):

Okay, 1. A 2. B

OpenStudy (anonymous):

Yea I literally just took the test a couple days ago xD

OpenStudy (theqube1):

So do you have all the answers? haha

OpenStudy (anonymous):

Yea I do. Can you do the explanation ones on your own? Here are the Multiple choice ones: 3. A 4. C 5. D 6. A 7. A 8. D 9. C 10. A

OpenStudy (theqube1):

Yeah I can try, thanks, I will let you know how I do in them lol

OpenStudy (anonymous):

I can help you with them if you want.

OpenStudy (anonymous):

I got all of them 100% right.

OpenStudy (theqube1):

Alright as soon as I gt to them. 1 moment

OpenStudy (anonymous):

Kayy

OpenStudy (theqube1):

Alright i'm here XD

OpenStudy (theqube1):

Here is the first opened ended question

OpenStudy (anonymous):

Okay! So, the area of a parallelogram is the same as an area of a rectangle: \[A = bh\]

OpenStudy (anonymous):

So, using that information, can you find the area of the parallelogram?

OpenStudy (theqube1):

Is it 1330?

OpenStudy (anonymous):

Yes! :D

OpenStudy (theqube1):

Yay!! Okay Next one?

OpenStudy (anonymous):

It's the area of a triangle right?

OpenStudy (theqube1):

yeah

OpenStudy (theqube1):

I got no idea how to do these

OpenStudy (anonymous):

\[A = 1/2 bh\] is the equation for the area of a triangle. simply plug those numbers in (and explain what you're doing obviously) and then you get your answer!

OpenStudy (theqube1):

How do I find base and height with that particular triangle?

OpenStudy (theqube1):

i'm thinking its 13.8?

OpenStudy (anonymous):

Sorry, back. Yep, it's 13.8!

OpenStudy (theqube1):

So what are the weird dotted lines for?

OpenStudy (anonymous):

that's the height of the triangle.

OpenStudy (anonymous):

It's there so you know what the height is and so it's less confusing.

OpenStudy (theqube1):

Ohhh okay, and area of the kit is the answer 60?

OpenStudy (anonymous):

Yep! You're progressing well :)

OpenStudy (theqube1):

thank ya :) So now this one I am not comprehending as much

OpenStudy (anonymous):

\[A = 1/2ap\] That's the equation for finding the area of regular polygons. They give you the apothem (a) and the side. All you have to do is multiply the side by 6 to get the perimeter, then plug those numbers in. :)

OpenStudy (theqube1):

is it 374.12?

OpenStudy (anonymous):

I got 374.4 ...

OpenStudy (anonymous):

I think it's close enough xD

OpenStudy (theqube1):

yeah I believe so haha

OpenStudy (anonymous):

Aiight, Next question, lemme at it xD

OpenStudy (theqube1):

alright here it is

OpenStudy (anonymous):

There's a theorem that I'm going to use with this problem: the Area of a Triangle Given SAS. THe theorem states that the area of any triangle is 1/2 the product of the lengths of two sides and the sine of the included angle. You'll find it in your textbook. Here's the equation: \[\Delta ABC = 1/2bc(\sin A)\]

OpenStudy (anonymous):

That's the area of a triangle given Side, Angle and Side. SAS.

OpenStudy (anonymous):

So, to fill those numbers in... \[\Delta ABC = \left( 533 \right) \left( 525 \right) \left( \sin 53° \right)\]

OpenStudy (anonymous):

You can use a setting on your calculator (I believe scientific) to get sin, cos, tan, etc.

OpenStudy (theqube1):

I got 62.6312° How farr off am I?

OpenStudy (anonymous):

We're not looking for an angle, remember, it's area. And I also made a mistake. you should multiply the numbers together and then multiply it by half.

OpenStudy (theqube1):

so then 111739 ft^2?

OpenStudy (anonymous):

Yep! Nice! :D

OpenStudy (theqube1):

Yay

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Alright, we're down to the last question. :)

OpenStudy (theqube1):

yes indeed

OpenStudy (anonymous):

Alright. We need to find the diameter of the ball. you can use the equation for circumference: \[C = πd\] OR \[C = 2πr\] Switch it around like so: \[R = C \div 2π \] to find the radius :)

OpenStudy (anonymous):

Let me know when you got that done. :)

OpenStudy (theqube1):

will do.

OpenStudy (theqube1):

9.55?

OpenStudy (anonymous):

Yep! I estimated to 9.6, to make things easier. now, since the hoop's diameter is 18 we can divide the difference of 18 and 9.6 by 2 to get the distance between the hoop and the basketball!

OpenStudy (theqube1):

2 inches then?

OpenStudy (anonymous):

18 - 9.6 is 8.4, 8.4 / 2 is 4.2.

OpenStudy (theqube1):

Deeeerrrrrppppp I read that as 18 divided by 9.6 = 2 lol

OpenStudy (anonymous):

xDDD

OpenStudy (anonymous):

And we're done! You will guaranteed get a 100 on that unit test :)

OpenStudy (theqube1):

Thank you a billion :D

OpenStudy (anonymous):

No problem! :D

OpenStudy (theqube1):

now I gotta portfolio I think I gotta do so lets see what this hell on earth is

OpenStudy (theqube1):

-.- I hate math

OpenStudy (anonymous):

Oh that...I did a modified version, my teacher put it on the message board...I don't know if your teacher did the same... :/

OpenStudy (theqube1):

WHo is your teacher?

OpenStudy (anonymous):

Same =_=

OpenStudy (anonymous):

Mr. Trautmann

OpenStudy (theqube1):

Nope I have Mr. Sorrentino

OpenStudy (theqube1):

Not even close o.o We have this program called geogebra that we use to make geometrical figures on and take the screenshots on then we send the teacher the screen shots and like the questions that go with them, it looks like this

OpenStudy (anonymous):

Ah, GeoGebra. Most confusing, I must say. Yea, this time, I can't help. I'm so so so sorry! D:

OpenStudy (anonymous):

I gotta go. It was so nice helping you! I will talk to you tomorrow hopefully :) Bye!

OpenStudy (theqube1):

Don't worry usually there are step by steps to do it, and Thank you again, and yeam maybe tomorrow my friend :)